Bone has a young's modulus of about 18 times 10^9
Answer #1

since

\sigma =\frac{F}{A}=Y\frac{\Delta L}{ L_{o}}

=>\Delta L=\left ( \frac{F}{A} \right )\frac{L_{o}}{Y}

=>\Delta L=\left (163\times 10^{6})\times \left ( 0.48/9\times 10^{9} \right )=8.7mm